OpenStudy (anonymous):

-4/15 x9 would it be -36/15? explain

OpenStudy (jdoe0001):

\(\bf -\cfrac{4}{15}\times 9\implies -\cfrac{4}{15}\times\cfrac{9}{1}\implies \cfrac{-4\times 9}{15\times 1}\implies -\cfrac{36}{15}\)